The area of a rectangle is 6 yard-square. The length of the rectangle is one more than twice the width. Find the dimensions of the rectangle.
Answer by jorel1380(3719)

w(2w+1)=6
2w2+w=6
2w2+w-6=0
(2w-3)(w+2)=0
w=3/2 or -2
Throwing out the negative answer, we get the width to be 3/2yds., and the length to be 4 yds..